monkey69:

Choose the system of equations that matches the following graph: (5 points) a 3x + 7y = 7 5x + 7y = −63 b 3x + 7y = 7 5x − 7y = −63 c 3x − 7y = 7 5x + 7y = −63 d 3x − 7y = 7 5x − 7y = −63

ramen:

Can you please upload the following graph so we can compare the system of equations? Thanks

monkey69:

does that work

ramen:

Yes it does. Thanks. Now for the problem, are you able to use a graphing calculator? It will speed things up. Try using this one: https://www.desmos.com/calculator

monkey69:

i cant

monkey69:

i dont know how to use that

ramen:

You simply just type in the equation you have into the text box on the left hand side. Here I'll show you a picture.

ramen:

1 attachment
monkey69:

oh

monkey69:

i cant get it

ramen:

Wdym you can't get it?

monkey69:

like i dont know how to do it

ramen:

Ok I will just explain how to do this problem without the calculator then. Understood?

monkey69:

okay

monkey69:

are you there

ramen:

For A and B for the equation 3x+7y=7, we have to convert this to standard form y=mx+b. To do this, we simply just isolate y by itself on one side. So, we subtract the 3x from the left side and put it on the right side making it 7y=-3x+7. We are almost done. Now we just need to divide everything in the equation by 7 to get rid of the 7 by the y. This makes the equation y=-3/7x + 1. Understand?

monkey69:

yes

ramen:

Btw, we have to convert each of these equations on the problem into standard form so we can graph them without a calculator. Once we do so, we can manually graph each equation to find your answer.

monkey69:

yeah

monkey69:

i think is c

monkey69:

hello

ramen:

Nope not quite, the equation 3x − 7y = 7 is negative and according to the graph in the problem that you compare it to, the red line is positive, while in C, 3x − 7y = 7 is negative.